OpenStudy (anonymous):

The midpoint of line UV is (5, -11) The coordinates of one endpoint are U(3, 5) Find the cordinated of endpoint V. (((PLEASE HELP IVE BEEN STUCK ON THIS FOR 20 MINS)))

OpenStudy (anonymous):

@AlexandervonHumboldt2 @I_Luv_Skittles @Defiance @freckles

OpenStudy (anonymous):

explain it too please.

OpenStudy (anonymous):

will medal and fan.

OpenStudy (freckles):

so the midpoint we need to think averages the average of two numbers is the sum of those numbers divided by 2 the x coordinate of the midpoint is equal to the sum of the x-coordinates of the endpoint divided by 2 the y coordinate of the midpoint is equal to the sum of the y-coordinates of the endpoint divided by 2

OpenStudy (freckles):

\[5=\frac{3+x}{2} \\ -11=\frac{5+y}{2}\]

OpenStudy (defiance):

Would help but seems freckles has it ;).

OpenStudy (freckles):

solve the equations

OpenStudy (freckles):

I forgot to make the word endpoint plural oh well

OpenStudy (anonymous):

what is x??? how do I do 3+x if I dont know what x is where did you get x from??

OpenStudy (freckles):

did you read the descriptions above?

OpenStudy (freckles):

the "x-coordinate of the midpoint" is equal to the "sum of the x-coordinates of the endpoints" "divided by 2"

OpenStudy (freckles):

V is a point in the form (x,y)

OpenStudy (freckles):

the x-coordinate of V is x the y-coordinate of V is y since I chose V to be the point (x,y)

OpenStudy (anonymous):

I get that now but I dont understand the divison part??? can you just write everything out for me to read please

OpenStudy (freckles):

I'm sorry I thought I did... Can you tell me what you don't understand? Like the division? To find the average of two numbers... You do the sum of those numbers and then you divide 2.

OpenStudy (anonymous):

theyre numbers on a graph tho so I dont get it

OpenStudy (freckles):

So if I told you the average of x and 3 is 5 ? This doesn't make sense how I got this: \[\frac{x+3}{2}=5?\]

OpenStudy (anonymous):

Youre saying the answer is 5?

OpenStudy (anonymous):

Im really bad at this kinda stuff

OpenStudy (freckles):

no you are looking for V and V is a point on the cartesian plane

OpenStudy (freckles):

you are looking for the point (x,y)

OpenStudy (anonymous):

so the x cord is 5?

OpenStudy (freckles):

you are given by your question the following to find x and y: The average of x and 3 is 5. The average of y and 5 is -11.

OpenStudy (freckles):

how did you x is 5?

OpenStudy (anonymous):

so its (5, -11) ?

OpenStudy (freckles):

the midpoint is not what we are looking for

OpenStudy (freckles):

I already said we are looking for V , one of the endpoints which I called (x,y)

OpenStudy (anonymous):

this is taking so long.. can you just explain everything and help me with the answer

OpenStudy (freckles):

I'm trying too... so what do you not understand about the following equations: \[5=\frac{3+x}{2} \\ -11=\frac{5+y}{2}\]

OpenStudy (freckles):

all i did there was apply the average thing I mentioned

OpenStudy (anonymous):

how do I add 3 to x and 5 to y..

OpenStudy (freckles):

because that is all the midpoint is is averaging

OpenStudy (freckles):

you solve for x and y

OpenStudy (freckles):

multiply both sides by 2

OpenStudy (anonymous):

so its 3 and 12.5?

OpenStudy (freckles):

don't understand how you got that did you try multiplying both sides of the equation(s) by 2?

OpenStudy (anonymous):

I did then divided them

OpenStudy (freckles):

\[5=\frac{3+x}{2} \\ \text{ multiply both sides by 2 } \\ 10=3+x \\ \text{ now isolate } x\]

OpenStudy (anonymous):

can you tell me the answer I dont get it my teacher will help me tomorrow

OpenStudy (freckles):

so you don't want to try to subtract 3 on both sides to isolate x?

OpenStudy (anonymous):

it would just be x over 2 ....

OpenStudy (freckles):

you do understand we had 5=(3+x)/2 then I asked you to multiply 2 on both sides giving us 10=3+x the last step to find the x-coordinate of v is to subtract 3 on both sides 10-3=x why are you doing something different from what I suggest?

OpenStudy (freckles):

now try solving for y coordinate of V -11=(5+y)/2

OpenStudy (freckles):

it is the same kind of steps we went through to solve for x

OpenStudy (anonymous):

my computer stopped working but can you just tell me it

OpenStudy (anonymous):

Ive spent over an hour on one problem...

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

ANYONE PLEASE :( IVE SPENT AN HOUR+ ON THIS QUESTION.... @*louisa* @perii224 @AlexandervonHumboldt2 @mrqueest

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

please...

Vocaloid (vocaloid):

start with the x-coordinates (x+3)/2 = 5 solve for x

OpenStudy (anonymous):

can you just give me the answer :/ i really dont get it my teacher will help me tomorrow. Ive spent over an hour on this...

Vocaloid (vocaloid):

start by multiplying both sides by 2 (x+3)/2 * 2 = ?

OpenStudy (anonymous):

4

Vocaloid (vocaloid):

nope, not quite... hint: A/2*2 = A (x+3)/2*2 = ?

OpenStudy (anonymous):

2*2 is 4

Vocaloid (vocaloid):

does this make it more clear?

OpenStudy (anonymous):

I DONT KNOW WHAT THIS MEANS AHHHHHH

Vocaloid (vocaloid):

just follow the pattern

OpenStudy (anonymous):

um I dont know what youre talking about can you just tell me it or Ill get it wrong I dont have the time to spend any more time on this one question...

OpenStudy (anonymous):

@Vocaloid